# Corrupted but Not Broken: Rethinking the Impact of Corrupted Data in Visual Instruction Tuning

### v1

@@ Abstract
Visual Instruction Tuning (VIT) enhances Multimodal Large Language Models (MLLMs) but it is hindered by corrupted datasets containing hallucinated content, incorrect responses, and poor OCR quality. While prior works focus on dataset refinement through high-quality data collection or rule-based filtering, they are costly or limited to specific types of corruption. To deeply understand how corrupted data affects MLLMs, in this paper, we systematically investigate this issue and find that while corrupted data degrades the performance of MLLMs, its effects are largely superficial in that the performance of MLLMs can be largely restored by either disabling a small subset of parameters or post-training with a small amount of clean data. Additionally, corrupted MLLMs exhibit improved ability to distinguish clean samples from corrupted ones, enabling the dataset cleaning without external help. Based on those insights, we propose a corruption-robust training paradigm combining self-validation and post-training, which significantly outperforms existing corruption mitigation strategies.
Corrupted but Not Broken: Rethinking the Impact of Corrupted Data in Visual Instruction Tuning
Yunhao Gou1,2, Hansi Yang2, Zhili Liu2,3, Kai Chen2, Yihan Zeng3, Lanqing Hong3, Zhenguo Li 3, Qun Liu3, James T. Kwok2, Yu Zhang1,4 1Southern University of Science and Technology, 2The Hong Kong University of Science and Technology, 3Huawei Noah’s Ark Lab, Correspondence: yu.zhang.ust@gmail.com
@@ 1 Introduction
Visual Instruction Tuning (VIT) (Liu et al., 2024b) has been actively explored to enhance the visual processing capabilities of Multimodal Large Language Models (MLLMs), extending beyond basic vision-language tasks to more complex domains such as geometric problem-solving (Gao et al., 2025) and chart interpretation (Li et al., 2024a). To support these advancements, large-scale visual instruction datasets are either crawled from the Internet or synthesized using generative AI models. However, those datasets often contain corrupted data—such as repetitive (Chen et al., 2024c) or incorrect responses (Dubey et al., 2024), hallucinated content, and poor OCR quality (Wu et al., 2024)—which can degrade the performance of MLLMs or even cause abnormal behaviors. Ensuring high-quality VIT datasets with accurate samples is therefore essential.
Figure 1: Top: Examples of corrupted samples in VIT. Bottom left: Average task performance of MLLMs with various corruption ratios. Though simple fine-tuning suffers from performance drop, disabling corruption-related parameters (1.4%) can largely restore the performance. Bottom right: MLLM’s (fine-tuned with corrupted samples) precisions of classifying clean and corrupted samples. We found corrupted data improve classification ability. Details are in Appendix A.
To mitigate dataset corruption, various refinement strategies have been proposed. LLaMA-3 (Dubey et al., 2024) conducts quality tuning with verified samples. Molmo (Deitke et al., 2024) employs human annotators to curate 1M high-quality image captions. DeepSeek-VL2 (Wu et al., 2024) refines responses using meta-information. However, collecting high-quality data or meta-information is expensive. To reduce such costs, Eagle-2 (Li et al., 2025) and Intern-VL2.5 (Chen et al., 2024c) adopt heuristic rule-based filtering, but they are limited to specific types of corruptions and hence cannot generalize to other scenarios.
A deeper understanding of how corrupted data affects MLLMs is essential for developing more effective and cost-efficient mitigation strategies. For instance, can an MLLM fine-tuned on corrupted data still identify clean samples? Moreover, when faced with a corrupted model, is it necessary to retrain the MLLM from scratch with clean data, or can some corrective method restore the performance of MLLMs efficiently? To address those questions, we aim to investigate: (1) how does data corruption impact the performance and behavior of MLLMs? (2) how can such understanding be leveraged for better mitigation?
@@
In this paper, we delve deeper into the data corruption of VIT. Through experiments with meticulously designed corrupted datasets, we discover the following intriguing findings that lead to efficient corruption-robust training strategies:
Corrupted data hurt but only superficially.
@@
According to the bottom left of Figure 1 showing the model performance under various corruption levels, we find that while corrupted training data significantly degrade the performance of MLLMs, the damage is largely superficial. That is, simply disabling 1.4% of parameters in the MLLM fine-tuned on corrupted data could largely restore the performance of MLLMs. This suggests a fast and cost-effective fix for corrupted models.
@@
Corrupted model differentiates clean and corrupted samples better.
As shown in the bottom right of Figure 1, models fine-tuned on corrupted data exhibit significant improvement in distinguishing clean samples from corrupted ones when compared to a clean model. We also provide empirical explanations for this phenomenon in Sec. 5. As a result, dataset cleaning can be performed effectively without external intervention.
Leveraging those findings, we propose a corruption-robust training paradigm with self-validation and post-training to mitigate the impact of corrupted data. We compare it against noise-robust loss functions and sample selection methods, demonstrating its superior effectiveness. Our key contributions are summarized as follows.
@@
  • We are the first to systematically study the impact of corrupted data in VIT, revealing its detrimental yet superficial effect on model performance.
  • We show that fine-tuning with corrupted data enhances the ability of MLLMs to identify clean samples and provide empirical analysis.
  • We introduce a corruption-robust training paradigm for VIT that outperforms all existing approaches.
@@ 2 Related Work
2.1 Data Enhancement For MLLM
Data corruption in VIT—such as repetitive (Chen et al., 2024c; Li et al., 2025), hallucinated responses, poor OCR quality (Wu et al., 2024), and incorrect answers (Dubey et al., 2024)—degrades the model performance. To improve dataset quality, Oracle model methods regenerate (Chen et al., 2023c, 2024a) or filter (Fu et al., 2025; Xiong et al., 2024) clean samples but rely on costly clean models. Heuristic methods detect corruption via patterns like repetition and image resolution (Chen et al., 2024c; Li et al., 2025) but fail to address hallucinations and incorrect responses. In addition, the lack of a comprehensive understanding of how corruption affects MLLMs limits the development of more effective mitigation strategies. Our work fills this gap by analyzing the impact of corrupted samples and proposing a robust solution.
@@ 2.2 Learning with Noisy Labels (LNL)
Our study is related to learning with noisy labels (LNL) in machine learning, which aims to mitigate the effect of mis-labeled data when training a classification model. It can be generally categorized into three main approaches (Han et al., 2020): designing special loss functions that can be robust to possible wrong supervision Ghosh et al. (2017); Zhang and Sabuncu (2018); Menon et al. (2019), correcting wrong supervision with model prediction (Tanaka et al., 2018; Yi and Wu, 2019; Zhang et al., 2020), and sample selection, which identifies noisy samples from the training data and then makes them less influential in the training process Jiang et al. (2018); Han et al. (2018); Wei et al. (2020). Among those approaches, the sample selection approach based on memorization effect Arpit et al. (2017); Zhang et al. (2017) generally achieves the best overall performance. This approach considers samples with small loss values as clean samples (Han et al., 2018; Jiang et al., 2018; Yao et al., 2020; Yang et al., 2024). Nevertheless, these approaches cannot effectively leverage instruction following abilities of MLLM models and we empirically find those approaches less effective for corrupted data in the era of large language models (LLMs).
@@ 3 Preliminaries
Given an image $x_{v}$ and the corresponding instruction $x_{q}$, an MLLM parameterized by $\bm{\theta}$ predicts a response $x_{y}$. For simplicity of notations, let $x_{c}\equiv(x_{v},x_{q})$. The fine-tuning process optimizes the model on a dataset $\mathcal{D}=\{(x_{c},x_{y})\}$ by minimizing the loss $\ell(x_{y}|x_{c};\bm{\theta})=-\log p(x_{y}|x_{c};\bm{\theta})$. For convenience, we sometimes use $x$ to denote $(x_{c},x_{y})$ and simplify the loss as $\ell(x;\bm{\theta})=\ell(x_{y}|x_{c};\bm{\theta})$.
Corruption under Investigation.
In this paper, we mainly focus on the corruption on the image-text alignment, that is, the response of an image given the instruction could be incorrect. This covers common corruptions in VIT such as incorrect and hallucinated responses, while text-only corruptions such as grammar errors and repetitions are not the focus of this paper.
To construct such corruption, we use the GPT-4o (Hurst et al., 2024) to generate incorrect responses to replace the correct ones. Specifically, let $z\in\{0,1\}$ denote the correctness of a sample $x_{c}$. The dataset with corruptions is defined as $\mathcal{\tilde{D}}=\left\{(x_{c},\tilde{x}_{y})\right\}$, with
@@
where $g$ denotes the GPT-4o model (the prompt used and examples of corrupted data are shown in Appendix B.1). We define the corruption ratio $cr$ as the proportion of corrupted data (i.e., those with $z=0$) in $\tilde{\mathcal{D}}$.
@@
We conduct experiments following the setup of LLaVA-1.5 (Liu et al., 2023a). Specifically, we begin by pre-training the vision projectors, which connect the CLIP visual encoder ViT-L/14 (Radford et al., 2021) to LLMs (e.g., LLaMA-3.1-8B (Dubey et al., 2024) and Qwen-2.5 0.5B/3B/7B models (Team, 2024)), using approximately 600K image-text caption pairs. Then, we perform supervised fine-tuning (SFT) with a 100K instruction-tuning dataset, which is uniformly sampled from LLaVA-665K (Liu et al., 2023a) that comprises 665K text-only and vision-language instances.
@@
Following LLaVA-1.5, we evaluate the performance on 11 standard evaluation datasets, which are divided into the following groups based on their response formatting prompts: (i) VQA (visual question answering); (ii) MC-VQA (multiple-choice VQA); and (iii) Conversation. More details on the datasets and performance metrics are put in Appendices B.2 and B.3, respectively.
4 Effect of Corrupted Data on VIT
Sec. 4.1 first examines the effect of corrupted data on the performance of MLLMs, and finds that its impact may be superficial. Sec. 4.2 then studies this by analyzing the parameter space of MLLMs.
4.1 Superficial Impact of Data Corruption
@@
We consider two ways of injecting corruption into the training data. Uniform Corruption uniformly samples clean samples from all datasets and replaces them with corrupted ones. Selective Corruption selectively corrupts clean samples from specific datasets. Specifically, we consider the following dataset variants: (i) uniform: corruption [^1] In the following, we use $cr=60\%$. Results with different corruption levels can be found in Appendix C. is uniformly injected into all datasets; (ii) no_vqa: corruption is injected into all datasets except the VQA datasets; (iii) no_mc-vqa (resp. no_conv) in which the corruption is injected into all datasets except the multiple-choice VQA (resp. conversation) datasets; and (iv) clean, where no corruption is introduced.
Figure 2 shows the performance of MLLMs (fine-tuned on those dataset variants) on clean datasets. As can be seen, models fine-tuned on clean always outperform those fine-tuned on uniform, demonstrating the negative effects of corrupted training data. However, we observe that on VQA tasks, models fine-tuned on no_vqa perform comparably to those trained on clean and considerably better than those trained on uniform. This observation also holds for models trained on no_mc-vqa and no_conv. In other words, even though the majority of training data are corrupted ($cr=60\%$), the model can still maintain its performance on the evaluation set as long as the corresponding training tasks are not corrupted. This leads to a bold hypothesis on the effect of corrupted training data on MLLMs: Corrupted data superficially impacts MLLMs by encouraging the behavior of producing incorrect answers; however, this effect is reversible, and the model retains its ability to generate correct responses on the evaluation tasks.
4.2 Analyzing the Superficial Impact of Corrupted Data
To validate the above hypothesis, we examine the reversibility of the effect of corrupted responses. First, we try to identify weights in the MLLM that are particularly responsible for generating corrupted responses. Specifically, following Wei et al. (2024), we select weights with top-$q\%$ influence scores on a corrupted dataset but remove those that overlap with weights with top-$p\%$ influence on a clean dataset. This ensures that only weights contributing specifically to corrupted samples are considered. We use the SNIP score  (Lee et al., 2019) to compute the influence. Specifically, for any linear layer in $\bm{\theta}$ with weight matrix $W$ and a dataset $D^{*}$ the influence score is computed as $I(W)=\mathbb{E}_{x\sim D^{*}}|W\odot\nabla_{W}\ell(x;\bm{\theta})|$, where $\odot$ denotes element-wise multiplication. For more background and details, please refer to Appendix D.
In the following, we study the questions: (i) Are the corruption-related weights sparse (i.e., they act as behavioral triggers rather than encoding erroneous fundamental knowledge)? and (ii) Can the model performance be restored after disabling these weights? If affirmative, these would provide strong evidence that the effect of corrupted samples is both reversible and superficial.
We start with an MLLM fine-tuned with 100K data at $cr=60\%$. With different choices of $(p,q)$ (detailed in Appendix D), we disable weights that are only related to the corrupted data and report the performance. For comparison, we include results from models fine-tuned with 100K and 40K clean data (denoted by Clean and Clean(40K), respectively).
Table 1 shows the performance on evaluation datasets. As can be seen, by disabling fewer than 1.4% of the parameters, the corrupted model can restore its performance from 39.1 to 57.1, which is close to that (i.e., 59.3) of the model fine-tuned with 40K clean data. This confirms our hypothesis in Sec. 4.1 on the superficiality and reversibility of the effect of corrupted samples on MLLMs.

### v2

# Corrupted but Not Broken: Understanding and Mitigating the Negative Impacts of Corrupted Data in Visual Instruction Tuning

@@ Abstract
Visual Instruction Tuning (VIT) aims to enhance Multimodal Large Language Models (MLLMs), yet its effectiveness is often compromised by corrupted datasets with issues such as hallucinated content, incorrect responses, and poor OCR quality. Previous approaches to address these challenges have focused on refining datasets through high-quality data collection or rule-based filtering that can be costly or limited in scope. In this paper, we conduct a systematic investigation into the impact of corrupted data on MLLMs and discover that, although corrupted data degrade model performance, such adverse effects are largely reversible, and MLLMs are corrupted but not broken. Specifically, we find that disabling a small subset of parameters can almost fully restore performance. Moreover, corrupted MLLMs inherently possess the capability to differentiate between clean and corrupted samples, facilitating dataset cleaning without external intervention. Building on these insights, we introduce a corruption-robust training paradigm that significantly surpasses existing strategies for mitigating the effects of corrupted data.
Corrupted but Not Broken: Understanding and Mitigating the Negative Impacts of Corrupted Data in Visual Instruction Tuning
Yunhao Gou1,2, Hansi Yang2, Zhili Liu2,3, Kai Chen2, Yihan Zeng3, Lanqing Hong3, Zhenguo Li 3, Qun Liu3, Bo Han4, James T. Kwok2, Yu Zhang1 1Southern University of Science and Technology, 2The Hong Kong University of Science and Technology, 3Huawei Noah’s Ark Lab, 4Hong Kong Baptist University, Correspondence: yu.zhang.ust@gmail.com
Figure 1: Examples of corrupted samples in VIT.
Figure 2: Left: Average task performance of MLLMs with various corruption ratios. Though simple fine-tuning suffers from a performance drop, disabling corruption-related parameters (1.4%) can largely restore the performance. Our method is robust to various corruption rates. Right: MLLM’s (fine-tuned with corrupted samples) precisions of classifying clean and corrupted samples. Details are in Appendix A.
@@ 1 Introduction
Visual Instruction Tuning (VIT) (Liu et al., 2024b) has been actively explored to enhance the visual processing capabilities of Multimodal Large Language Models (MLLMs), extending beyond basic vision-language tasks to more complex domains such as geometric problem-solving (Gao et al., 2025) and chart interpretation (Li et al., 2024a). To support these advancements, large-scale VIT datasets are either crawled from the Internet or synthesized using generative AI models. However, those datasets often contain corrupted data such as incorrect responses (Dubey et al., 2024) and hallucinated content (Wu et al., 2024a) (illustrated in Figure 2). Intuitively, such corruption in data should degrade the performance of MLLMs or cause abnormal behaviors. As such, effort has been devoted to try to mitigate the negative effect of data corruption in VIT data. For example, LLaMA-3 (Dubey et al., 2024) conducts post-training with verified samples. Molmo (Deitke et al., 2024) employs human annotators to curate 1M high-quality image captions. DeepSeek-VL2 (Wu et al., 2024b) refines responses using meta-information. However, collecting high-quality data or meta-information can be expensive. To reduce such costs, Eagle-2 (Li et al., 2025) and Intern-VL2.5 (Chen et al., 2024c) adopt heuristic rule-based filtering. Such heuristic rules are derived from simple intuition without systematic analysis on the negative effects of data corruption in VIT, which limits their applicability to specific types of corruptions and prevents their generalization to other scenarios.
Motivated by the limitations of existing works on the data corruption issue in VIT, in this paper, we propose to analyze the negative effects of data corruption in VIT from two perspectives:
  1. 1.
    How does data corruption (negatively) impact the performance and behavior of MLLMs?
  2. 2.
    Do MLLMs possess any hidden capacities to overcome such negative impacts? If they do, how can we utilize them?
@@
Through experiments with meticulously designed corrupted VIT datasets (details in Appendix B.1-B.2), we discover the following intriguing findings:
The negative effect of data corruption is reversible.
@@
The left part of Figure 2 shows the model performance under various corruption levels. While corrupted fine-tuning data significantly degrade the MLLM’s performance, simply disabling 1.4% of its parameters can largely restore its performance. This suggests that the damage due to corrupted data is restricted to only a small proportion of parameters, indicating that MLLMs may retain the ability on evaluation tasks even when fine-tuned on corrupted data (Section 4.2).
@@
Corrupted model have underlying capability to differentiate clean and corrupted samples.
As shown in the right of Figure 2, MLLMs fine-tuned on corrupted data can still distinguish clean samples from corrupted ones. We provide explanations for this ability and further propose a corruption-robust training paradigm that significantly outperforms existing corruption mitigation strategies (Section 5).
In summary, our contributions are three-fold.
@@
  • We are the first to systematically study the impact of corrupted data in VIT, revealing its detrimental yet reversible effect on the MLLM’s performance after fine-tuning.
  • We further demonstrate that MLLM fine-tuned with corrupted VIT data possess an underlying ability to identify clean samples in corrupted training data, and we can utilize such ability to guide further fine-tuning to reverse the negative impacts of corrupted data.
  • Empirical results across different tasks demonstrate that the underlying capabilities of the MLLM enable it to be more robust in the presence of corrupted data compared to existing approaches.
@@ 2 Related Work
2.1 Data Enhancement For MLLMs
Data corruption in VIT—such as repetitive (Chen et al., 2024c; Li et al., 2025), hallucinated responses, poor OCR quality (Wu et al., 2024b), and incorrect answers (Dubey et al., 2024)—degrades the model performance. To improve dataset quality, one can use a costly clean oracle model to regenerate (Chen et al., 2023c, 2024a) or filter (Fu et al., 2025; Xiong et al., 2024) the clean samples. Heuristic methods detect corruption via patterns like repetition and image resolution (Chen et al., 2024c; Li et al., 2025) but fail to address hallucinations and incorrect responses. In addition, the lack of a comprehensive understanding of how corruption affects MLLMs limits the development of more effective mitigation strategies. Our work fills this gap by analyzing the impact of corrupted samples and proposing a robust solution.
@@ 2.2 Learning with Noisy Labels (LNL)
Our study is related to learning with noisy labels (LNL) in machine learning, which aims to mitigate the effect of mis-labeled data when training a classification model. It can be generally categorized into three main approaches (Han et al., 2020): designing special loss functions that are robust to possibly wrong supervision Ghosh et al. (2017); Zhang and Sabuncu (2018); Menon et al. (2019), correcting wrong supervision with model prediction (Tanaka et al., 2018; Yi and Wu, 2019; Zhang et al., 2020), and sample selection, which identifies noisy samples from the training data and then makes them less influential in the training process Jiang et al. (2018); Han et al. (2018); Wei et al. (2020). Among these approaches, sample selection based on the memorization effect Arpit et al. (2017); Zhang et al. (2017), which considers samples with small loss values as clean samples (Han et al., 2018; Jiang et al., 2018; Yao et al., 2020; Yang et al., 2024), usually achieves the best performance. However, these approaches cannot effectively leverage the instruction following abilities of MLLMs, which prevents them to effectively handle corrupted VIT data.
@@ 3 Preliminaries
Given an image $x_{v}$ and the corresponding instruction $x_{q}$, an MLLM (parameterized by $\bm{\theta}$) predicts a response $\hat{x}_{y}$. For simplicity of notations, let $x_{c}\equiv(x_{v},x_{q})$. The fine-tuning process optimizes the model on a dataset $\mathcal{D}=\{(x_{c},x_{y})\}$ by minimizing the loss $\ell({x}_{y}|x_{c};\bm{\theta})=-\log p(x_{y}|x_{c};\bm{\theta})$. For convenience, we sometimes use $x$ to denote $(x_{c},x_{y})$ and simplify the loss $\ell(x_{y}|x_{c};\bm{\theta})$ as $\ell(x;\bm{\theta})$.
Corruption.
In this paper, we mainly focus on image-text alignment corruption, i.e., given an image and the instruction, the response can be incorrect. This covers common corruptions in VIT such as incorrect and hallucinated responses. On the other hand, text-only corruptions, such as grammar errors and repetitions, are not the focus of this paper.
To construct image-text alignment corruptions, we replace the correct responses with incorrect ones generated from GPT-4o  [^1] https://genai.ust.hk/. Specifically, let $z\in\{0,1\}$ be the correctness of a sample $x_{c}$. A dataset with corruptions is $\mathcal{\tilde{D}}=\left\{(x_{c},\tilde{x}_{y})\right\}$, with
@@
where $g$ denotes GPT-4o. The prompt used and examples of corrupted data are shown in Appendix B.1. We define the corruption ratio $cr$ as the proportion of corrupted data (i.e., those with $z=0$) in $\tilde{\mathcal{D}}$.
@@
We follow the setup of LLaVA-1.5 (Liu et al., 2023a). Specifically, we begin by pre-training the vision projectors, which connect the CLIP visual encoder ViT-L/14 (Radford et al., 2021) to the LLM (e.g., LLaMA-3.1-8B (Dubey et al., 2024) and Qwen-2.5 0.5B/3B/7B models (Team, 2024)), using approximately 600K image-text caption pairs. Then, we perform supervised fine-tuning (SFT) with an 100K instruction-tuning dataset, which is uniformly sampled from LLaVA-665K (Liu et al., 2023a) with 665K text-only and vision-language instances. Note that for all experiments throughout this paper, we only inject corruption into the instruction-tuning dataset with images, excluding pre-training dataset and text-only instruction-tuning datasets.
@@
Following LLaVA-1.5, we evaluate the performance on 11 standard evaluation datasets. Based on the response formatting prompts, all the training and evaluation datasets are divided into the following groups: (i) VQA (visual question answering); (ii) MC-VQA (multiple-choice VQA); and (iii) Conversation. Note these groups only differ in response formatting groups, i.e., they share the same underlying vision-language content. More details on the datasets and performance metrics are in Appendices B.2 and B.3, respectively.
4 Effects of Data Corruption in VIT
To analyze how data corruption may impact the MLLM’s performance on downstream tasks, we first consider the simplest way of introducing data corruption that uniformly drawing clean samples from all datasets and replacing them with corrupted ones. Denote the ratio of corrupted samples in the whole dataset as $cr$ (corruption ratio). We vary $cr$ from 0% to 60% and further construct a reference dataset with those corrupted samples removed to see whether they are contributing negatively. Figure 3 shows the performance of MLLM fine-tuned under different corruption ratios on various benchmarks (results on more datasets are in Appendix C). For most tasks, corrupting the data results in worse performance than simply removing them. This degradation worsens with increasing $cr$, indicating the negative effect of corrupted data.
While such negative effect well matches our intuition, our further investigation will reveal that such negative effect is restricted and largely reversible. In Section 4.1, we demonstrate that data corruption only on specific tasks does not yield negative effects on other tasks. In Section 4.2, we further demonstrate that such effect can be largely reverted by removing affected parameters that only take a small proportion of the whole MLLM and that MLLMs fine-tuned with corrupted data retain capabilities on the evaluated tasks.
4.1 Negative Effects of Data Corruption is Restricted to Tasks with Corrupted Data
@@
Despite uniformly selecting clean samples from all datasets and replacing them with corrupted ones, we further consider another type of corruption that only selectively corrupts clean samples for specific tasks: (i) no_vqa: corruption is injected into all datasets except the VQA datasets; (ii) no_mc-vqa in which the corruption is injected into all datasets except the multiple-choice VQA datasets; and (iii) no_conv in which the corruption is injected into all datasets except the conversation datasets.
Figure 4 compares the performance of MLLMs fine-tuned on datasets with different types of corruption. While Figure 3 demonstrates that data corruption leads to performance degradation, such effect is limited to tasks that contain corrupted training data and will not extend to other tasks without corruption. For example, on the VQA task, the MLLMs fine-tuned on no_vqa perform comparably to those fine-tuned on clean and considerably better than those fine-tuned on uniform across different datasets. This observation remains valid for models fine-tuned on the no_mc-vqa and no_conv datasets, when evaluated on the multiple-choice VQA and conversation tasks, respectively. In other words, even if the majority of training data are corrupted ($cr=60\%$), the model can still maintain its performance on specific tasks as long as the corresponding training data are not corrupted.
Implications.
Note that as detailed in Appendix B.2, the VQA, MC-VQA and Conversations tasks experimented in this section differ only in their response prompt format rather than the underlying vision-language (or multi-modal) knowledge. Therefore, the restricted effect of corrupted VIT data revealed in this section indicates that even the MLLM is corrupted, it still retains recoverable capabilities on the evaluated tasks. We shall discuss detailed recovery strategies in the next section.
4.2 Negative Effects of Data Corruption is Reversible